OpenStudy (anonymous):

Solve the system of equations: y + 4x = 8 and 5x + 2y = 13. Answer (-1, 4) (1, 4) (1, 12) (-1, 12)

OpenStudy (anonymous):

\[2y+8x=16\] \[(2y+8x)-(2y+5x)=16-13\] \[3x=3\] \[x=1\] \[2y+8=16\]\[y=4\]

OpenStudy (anonymous):

Thnx

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

Sure thing

OpenStudy (anonymous):

Solve the system of equations: 2x + 2y = 6 and 12 = 2y – x. Answer (2, 5) (-4, 5) (-2, 5) (-2, 7)

OpenStudy (anonymous):

What's the second equation?

OpenStudy (anonymous):

Solve the system of equations: 2x + 2y = 6 and 12 = 2y – x.

OpenStudy (anonymous):

\[2x+2y-2y=6-12\] \[2x=-6\] \[x=-3\]\[y=6\]

OpenStudy (anonymous):

its not one of the options though

OpenStudy (anonymous):

I can't see your second equation. All I see is 12=2y and then a whole bunch of ? marks

OpenStudy (anonymous):

Solve the system of equations: 2x + 2y = 6 and 12 = 2y – x. Answer (2, 5) (-4, 5) (-2, 5) (-2, 7)

OpenStudy (anonymous):

Perhaps you could type out the second equation

OpenStudy (anonymous):

solve the system of equations 2x+2y=6 and 12 = 2y - x a) (2,5) b) (-4,5) c) (-2,5) d) (-2,7)

OpenStudy (anonymous):

:) much better \[(2x+2y)-(2y-x)=-6\] \[2x-(-x)=2x+x=3x=-6\]\[x=-2\] \[12=2y-(-2)\] \[y=5\]

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Glad I could help

OpenStudy (anonymous):

i only have 2 more left. but i can ask someone else. if u dont want to. once again thnx!

OpenStudy (anonymous):

I don't mind

OpenStudy (anonymous):

do u want to?

OpenStudy (anonymous):

Sure, why not

OpenStudy (anonymous):

Solve the system of equations 10x + 7y = 49 and 10y -x = 70 a) No solution b) (0,7) c) (7,0) d) (-7, 1)

OpenStudy (anonymous):

x=0 y=7

OpenStudy (anonymous):

Last one, Solve the systems of equations: 2x + 9y = 0 and 3x +5y = 17 a) (9, -2) b) (-9,-2) c) (9,2) d) No solutions

OpenStudy (anonymous):

\[-2x/9=y\] \[3x+5(-2x/9)=17\] \[27x-10x=151\] \[17x=151\] \[x=9\] \[18+9y=0\] \[y=-2\]

OpenStudy (anonymous):

Thank u so much! i appreciate it!

OpenStudy (anonymous):

You're very welcome

OpenStudy (dan815):

@abb0t